find f(x) = x2 - 2x -2 find f' (x) using the long method


1
Expert's answer
2021-11-02T16:42:48-0400

"\displaystyle
f(x) = x^2 -2x -2\\
f'(x) = \lim _{h \to 0}\frac{|f(x+h)-f(x)|}{|h|} \\
\text{where $f(x+h) = (x+h)^2 -2(x+h) -2$ }\\
=x^2+2xh+h^2-2x-2h-2\\
\therefore \frac{f(x+h)-f(x)}{h}=2x+h-2\\
\therefore \lim_{h \to 0}\frac{|f(x+h)-f(x)|}{|h|}=2x-2"